OpenStudy (anonymous):

How to make m^13/m^8 a single exponent?

jimthompson5910 (jim_thompson5910):

When dividing variables with the same base, you subtract the exponents So 13 - 8 = 5 That 5 is the exponent for the base of m In other words, m^13/m^8 = m^5

OpenStudy (the_fizicx99):

The bases are the same so you would subtract 13 - 8 m^5